# A Survey of Model Architectures in Information Retrieval

### v2

@@ 6 IR with Pre-trained Transformers
The most effective application of transformers in IR involves full, deep interaction between query and document tokens. This architecture, known as a cross-encoder (Humeau et al., 2020), takes the concatenated sequence of $(\mathcal{Q},\mathcal{D})$ as a single input. Nogueira et al. (2019b) first employed this approach in their monoBERT model for reranking candidate passages from a first-stage retriever. The model outputs a relevance score $s$ via a linear layer on top of the final BERT representation, typically from a linear layer using the [CLS] token’s representation (Fig.˜3(a)). Conceptually similar to pre-Transformer interaction-based neural ranking models, this schema has proven effective across various pre-trained encoders (Zhang et al., 2021), as well as other transformer architectures architectures (§˜7). However, this cross-encoder schema faces two primary challenges: (1) the fixed context length of models like BERT (e.g., 512 tokens) makes processing long documents difficult, and (2) relying on a single token’s fixed-dimensional representation (e.g., 768-dimensional representation for BERT) may limit the model’s expressive power.
@@
Corresponding mitigations for these two challenges have been extensively investigated in literature. Chunk-and-aggregate approaches represent a practical solution to handle long documents that exceed BERT’s input constraints by decomposing the ranking problem into passage-level scoring followed by aggregation (Gao & Callan, 2022). The fundamental strategy involves splitting documents into fixed-length passages or sentences, applying BERT-based cross-encoders to score each query-passage pair independently, then combining these scores to produce a final document-level relevance score. Early work in this direction explored sentence-level aggregation, where BERT scores computed at the sentence level were shown to be effective for document ranking (MacAvaney et al., 2020; Yilmaz et al., 2019). The BERT-MaxP (Dai & Callan, 2019a) approach became particularly influential, where documents are split into fixed-length passages and the maximum passage score serves as the document score. Two primary aggregation strategies have emerged: (1) score-pooling and (2) representation aggregation. Score-pooling methods apply simple operations like maximum, sum, or first passage scores to combine passage-level relevance scores (Dai & Callan, 2019a). In contrast, representation aggregation methods address both the long-document problem and the single-vector expressiveness limitation. Instead of collapsing each passage’s signal into a single scalar score, these approaches gather the rich, low-dimensional [CLS] token representations from each passage. This collection of vectors forms a more comprehensive and expressive document-level feature set, which is then processed by additional neural networks (MacAvaney et al., 2020). Notable systems like PARADE (Li et al., 2020) employ CNNs and transformers for aggregation, while CEDR (MacAvaney et al., 2019) pioneered joint approaches that combine BERT outputs with existing neural IR models through averaging.
@@
While chunk-and-aggregate approaches successfully handle long documents, they fundamentally limit query-document interactions to the passage level, creating information bottlenecks where passage scores or low-dimensional representations constrain the model’s ability to capture document-wide relevance patterns. Hofstätter et al. (2021b) argue that this trade-off between scalability and interaction richness remains a defining characteristic of this approach.
Efficient Pre-computation Models: The Bi-Encoder Architecture.
@@
While cross-encoders offer state-of-the-art effectiveness, their computational cost — requiring a full transformer pass for every $(\mathcal{Q},\mathcal{D})$ pair — makes them infeasible for retrieval over large collections. This limitation motivated the development of bi-encoder architectures, which are conceptually similar to representation-based neural ranking models (§˜5). [^2] The term “bi-encoder” is also known as a two-tower architecture or an embedding model. We use “bi-encoder” to contrast with “cross-encoder”, which takes concatenated input. A bi-encoder uses a backbone network (typically a transformer) to encode the query $\mathcal{Q}$ and document $\mathcal{D}$ separately. The resulting dense vector representations are then used to compute a relevance score with a simple similarity function like dot product or cosine similarity (Xiong et al., 2020; Karpukhin et al., 2020; Gao et al., 2021b). The key advantage of this separation is efficiency: the entire document collection can be encoded into a vector index offline. At query time, retrieval becomes a fast approximate nearest neighbor (ANN) search problem (Johnson et al., 2019; Malkov & Yashunin, 2016) or search with an inverted index data structure (Zobel et al., 1998), avoiding costly neural network inference. A notable insight from Lin (2021) is that the bi-encoder framework provides a unifying lens for understanding diverse retrieval approaches. Dense retrieval models, learned sparse retrieval models, and traditional bag-of-words approaches like BM25 can all be viewed as parametric variations of this architecture, differing primarily along two dimensions: the representation basis (dense semantic vectors vs. sparse lexical vectors) and whether the representations are learned or hand-crafted. Existing methods based on this bi-encoder architecture vary primarily in their representation format (dense vs. sparse), pooling strategies, and training methodologies.
Learned Dense Retrieval.
@@
Dense retrieval models employ a standardized dual encoder architecture built on pre-trained transformer models, most common BERT in this section. The standard formulation uses separate BERT encoders for queries and documents, with a layer-normalized linear projection applied to the token representation: $\text{Encoder}(\cdot)=\text{Linear}\big{(}\text{BERT}(\cdot)\big{)}$. The encoder weights can be separate or shared between query and document sides.
@@
Relevance scoring in dense retrieval is performed through simple similarity functions, most commonly dot product or cosine similarity between query and document vectors. This design enables efficient approximate nearest neighbor (ANN) search over pre-computed document representations (Johnson et al., 2019), making dense retrieval practical for large-scale collections while maintaining the semantic understanding capabilities of transformer models.
@@
The success of this architecture stems from its ability to learn semantic representations that address the vocabulary mismatch problem inherent in traditional sparse retrieval methods (Lee et al., 2019; Karpukhin et al., 2020; Xiong et al., 2020). Dense retrieval models have demonstrated notable effectiveness improvements over BM25 baselines across various tasks including open-domain question answering and web search.
Learned Sparse Retrieval.
@@
Learned sparse retrieval (LSR, Fig.˜3(c)) employs the same bi-encoder architecture as dense retrieval but produces fundamentally different representations. [^3] We focus on the learned sparse retrieval under the bi-encoder formulation, which excludes works including learned document expansion, e.g., Nogueira et al. (2019a); Zbib et al. (2019), and document term reweighting, e.g.,DeepCT (Dai & Callan, 2019b). See Mallia et al. (2021); Basnet et al. (2024) for a unified narrative for learned sparse retrieval that includes these lines of works. While sharing the transformer backbone, sparse retrieval models encode queries and documents into high-dimensional sparse vectors whose dimensionality typically matches the vocabulary size of the underlying pre-trained model, often containing tens of thousands of dimensions. Each dimension corresponds to a specific vocabulary term, creating an interpretable representation where non-zero weights indicate term importance (Formal et al., 2021b; a; Nguyen et al., 2023).
@@
Three key architectural constraints distinguish sparse encoders from their dense counterparts. First, sparsity is enforced through explicit regularization techniques, ensuring most term weights remain zero to maintain efficiency (Formal et al., 2021b; Xu et al., 2025a). Second, all weights must be non-negative to maintain compatibility with traditional inverted index software designed for lexical search systems like BM25. Third, the high-dimensional vocabulary-aligned vectors enable direct integration with existing inverted index infrastructure, providing computational efficiency advantages over approximate nearest neighbor search required by dense methods.
@@
At a conceptual level, learned sparse retrieval can be viewed as a sophisticated evolution of traditional term weighting schemes, learning context-aware token importance scores from data rather than relying on heuristic formulas (Zamani et al., 2018; Dai & Callan, 2019b; Mallia et al., 2021; Yu et al., 2024a; Xu et al., 2025a). This approach inherits desirable properties from bag-of-words models such as exact term matching while leveraging the semantic understanding capabilities of pretrained transformers. Notable implementations include SPLADE (Formal et al., 2021a), DeepImpact (Mallia et al., 2021), and uniCOIL (Lin & Ma, 2021), which demonstrate that transformer-based sparse representations can achieve effectiveness comparable to dense retrieval while maintaining the efficiency benefits of inverted indexes.
Bridging the Gap: Advanced Interaction and Hybrid Models.
@@
Multi-Vector Representations.
@@
To re-introduce query-document interaction without the full cost of a cross-encoder, multi-vector models represent queries and documents using multiple vectors. Poly-Encoder (Humeau et al., 2020) computes a fixed number of vectors per query and aggregates them with softmax attention over document vectors. ME-BERT (Luan et al., 2021) represents documents with $m$ vectors and uses the maximum similarity between any query and document vector to estimate relevance. In line with this idea, ColBERT (Khattab & Zaharia, 2020; Santhanam et al., 2022; Hofstätter et al., 2022) represent each token in the query and document as a contextualized vector. It then performs a “late interaction” step where each query vector is compared against all document vectors via a MaxSim operator, and the final score is the sum of these maximum similarities. This late interaction scheme (Fig.˜3(d)) allows ColBERT for end-to-end training to achieve strong performance while still achieving efficient retrieval through a dedicated index structure. On the other hand, it also leads to drastically increased index size, which has been the focus in later studies (Santhanam et al., 2022; Hofstätter et al., 2022, inter alia).
Hybrid Retrieval.
@@
Another direction combines the strengths of different retrieval systems. A simple yet effective approach is ranklist fusion (e.g., Reciprocal Rank Fusion, Cormack et al., 2009), which merges ranked lists from sparse (e.g., BM25) and dense retrievers post-retrieval without architectural changes. More integrated models combine signals at a deeper level. COIL (Gao et al., 2021a) enhances traditional bag-of-words retrieval with semantic embeddings from a BERT encoder. uniCOIL (Lin & Ma, 2021) simplifies this by reducing the semantic embedding to a single dimension, effectively learning a term weight akin to LSR models like SPLADE (Formal et al., 2021b; a). A few works fall into the intersection of learned sparse retrieval and multi-vector representations. For example, SLIM (Li et al., 2023b) first maps each contextualized token vector to a sparse, high-dimensional lexical space before performing late interaction between these sparse token embeddings. SPLATE (Formal et al., 2024) takes an alternative approach to first encode contextualized token vectors, then map these token vectors to a sparse vocabulary space with a partially learned SPLADE module. Both models achieve performance improvement compared to learned sparse retrieval baselines such as SPLADE (Formal et al., 2021b; a).
@@
CEDR (MacAvaney et al., 2019) Reranking Cross-encoder BERT Contrastive Learning
@@
SLIM (Li et al., 2023b) LSR + Multi-vector Bi-encoder BERT Contrastive Learning
SLIM++ (Li et al., 2023b) LSR + Multi-vector Bi-encoder BERT Distillation
@@
Table 3: Summary of IR model architecture for passage retrieval and passage ranking based on pre-trained transformers. Dense Retrieval and LSR denote learned dense retrieval and learned sparse retrieval, respectively. DeepCT (Dai & Callan, 2019b) is trained without labeled training set. The “late interaction” mechanism introduced in (Khattab & Zaharia, 2020; Santhanam et al., 2022) can be considered a special case of multi-vector architecture. Contrastive Learning and in-batch negatives means listwise loss function is used.
Orthogonal Improvements and Analysis.
Beyond architectural innovations, performance can be enhanced through improvements to the underlying models and a deeper analysis of their behavior. We show a list of models and their corresponding architectures in Table˜3, a majority of which use BERT (Devlin et al., 2019) as the backbone, with exceptions using DistilBERT (Sanh, 2019), RoBERTa (Liu, 2019), and T5 (Raffel et al., 2020; Sanh et al., 2022; Mo et al., 2023; Chung et al., 2024).
@@
Instead of modifying the retrieval architecture, this line of work enhances the backbone language model itself through domain adaptation or continued pre-training, a proven strategy in NLP (Howard & Ruder, 2018; Gururangan et al., 2020). For instance, Lee et al. (2019) pre-train BERT with an Inverse-Cloze Task (Taylor, 1953) for better text representations. Condenser (Gao & Callan, 2021) proposes a dedicated pre-training architecture to "condense" text representations into the [CLS] token. COCO-DR (Yu et al., 2022b) extends Condenser by using a technique named Distributionally Robust Optimization to mitigate distribution shifts in dense retrieval. A line of works have also explored other pre-training objectives such as masked auto-encoders (Xiao et al., 2022; Wu et al., 2023b) and bag-of-words prediction (Ma et al., 2024a). Recent works (Wang et al., 2023a; Nussbaum et al., 2025; Yu et al., 2024b, inter alia) have employed a “middle-stage” training on large-scale unlabeled text pairs, between pretrained encoder models and supervised finetuning on labeled text pairs, and have demonstrated the corresponding performance improvement compared to the traditional two-stage pipeline. We refer readers to the original papers for details.
@@
A few works have attempted to interpret what transformer-based models learn, i.e., mechanistic interpretability (Elhage et al., 2021; Saphra & Wiegreffe, 2024). MacAvaney et al. (2022) showed that neural models rely less on exact match signals and instead encode rich semantic information. Ram et al. (2023) connected dense and sparse retrieval by projecting a dense retriever’s intermediate representations into the vocabulary space. Separately, other work focuses on designing systems that provide model-agnostic explanations (Rahimi et al., 2021; Yu et al., 2022a; Xu et al., 2024c) to satisfy desiderata like faithfulness (Jacovi & Goldberg, 2020; Xu et al., 2023b). As IR systems become integral to other applied ML domains, we believe it is important to study and design interpretable, truthful, and trustworthy IR models.
The architectural innovations discussed in this section highlight a mature research field dedicated to harnessing pre-trained transformers for information retrieval. The central theme has been the architectural trade-off between interaction depth and computational cost, giving rise to a spectrum of models from highly effective cross-encoder rerankers to efficient bi-encoder retrievers. By developing advanced representations — be they dense, sparse, or multi-vector — and hybridizing different approaches, the community has pushed the boundaries of the classic "retrieve-then-rank" paradigm. However, these models still primarily function as specialized components for representation and scoring. The next wave of innovation would come from models capable not just of understanding text, but of generating it, leading to the era of Large Language Models.

### v3

@@ 6.2 Deep Interaction Models: The Cross-Encoder for Reranking
The most effective application of transformers in IR involves full, deep interaction between query and document tokens. This architecture, known as a cross-encoder (Humeau et al., 2020), takes the concatenated sequence of $(\mathcal{Q},\mathcal{D})$ as a single input. Nogueira et al. (2019b) first employed this approach in their monoBERT model for reranking candidate passages from a first-stage retriever. The model outputs a relevance score $s$ via a linear layer on top of the final BERT representation, typically from a linear layer using the [CLS] token’s representation (Figure˜3(a)).
Conceptually similar to pre-Transformer interaction-based neural ranking models, this schema has proven effective across various pre-trained encoders (Zhang et al., 2021b), as well as other transformer architectures (Section˜7). However, this cross-encoder schema faces two primary challenges: (1) the fixed context length of models like BERT (e.g., 512 tokens) makes processing long documents difficult, and (2) relying on a single token’s fixed-dimensional representation (e.g., 768-dimensional representation for BERT) may limit the model’s expressive power.
Handling Long Documents.
@@
Corresponding mitigations for these two challenges have been extensively investigated in literature. Chunk-and-aggregate approaches represent a practical solution to handle long documents that exceed BERT’s input constraints by decomposing the ranking problem into passage-level scoring followed by aggregation (Gao & Callan, 2022). The fundamental strategy involves splitting documents into fixed-length passages or sentences, applying BERT-based cross-encoders to score each query-passage pair independently, then combining these scores to produce a final document-level relevance score. Early work in this direction explored sentence-level aggregation, where BERT scores computed at the sentence level were shown to be effective for document ranking (MacAvaney et al., 2020; Yilmaz et al., 2019). The BERT-MaxP (Dai & Callan, 2019a) approach became particularly influential, where documents are split into fixed-length passages and the maximum passage score serves as the document score.
Two primary aggregation strategies have emerged: (1) score-pooling and (2) representation aggregation. Score-pooling methods apply simple operations like maximum, sum, or first passage scores to combine passage-level relevance scores (Dai & Callan, 2019a). In contrast, representation aggregation methods address both the long-document problem and the single-vector expressiveness limitation. Instead of collapsing each passage’s signal into a single scalar score, these approaches gather the rich, low-dimensional [CLS] token representations from each passage. This collection of vectors forms a more comprehensive and expressive document-level feature set, which is then processed by additional neural networks (MacAvaney et al., 2020). Notable systems like PARADE (Li et al., 2023a) employ CNNs and transformers for aggregation, while CEDR (MacAvaney et al., 2019b) pioneered joint approaches that combine BERT outputs with existing neural IR models through averaging.
@@
While chunk-and-aggregate approaches successfully handle long documents, they fundamentally limit query-document interactions to the passage level, creating information bottlenecks where passage scores or low-dimensional representations constrain the model’s ability to capture document-wide relevance patterns. Hofstätter et al. (2021b) argue that this tradeoff between scalability and interaction richness remains a defining characteristic of this approach.
6.3 Efficient Pre-Computation Models: The Bi-Encoder Architecture
While cross-encoders offer state-of-the-art effectiveness, their computational cost — requiring a full transformer pass for every $(\mathcal{Q},\mathcal{D})$ pair — makes them infeasible for retrieval over large collections. This limitation motivated the development of bi-encoder architectures, which are conceptually similar to representation-based neural ranking models (Section˜5). [^2] The term “bi-encoder” is also known as a two-tower architecture or an embedding model. We use “bi-encoder” to contrast with “cross-encoder”, which takes concatenated input.
@@ 6.3 Efficient Pre-Computation Models: The Bi-Encoder Architecture
A bi-encoder uses a backbone network (typically a transformer) to encode the query $\mathcal{Q}$ and document $\mathcal{D}$ separately. The resulting dense vector representations are then used to compute a relevance score with a simple similarity function like dot product or cosine similarity (Xiong et al., 2020; Karpukhin et al., 2020; Gao et al., 2021b). The key advantage of this separation is efficiency: the entire document collection can be encoded into a vector index offline. At query time, retrieval becomes a fast approximate nearest neighbor (ANN) search problem (Johnson et al., 2019; Malkov & Yashunin, 2018) or search with an inverted index data structure (Zobel et al., 1998), avoiding costly neural network inference.
A notable insight from Lin (2021) is that the bi-encoder framework provides a unifying lens for understanding diverse retrieval approaches. Dense retrieval models, learned sparse retrieval models, and traditional bag-of-words approaches such as BM25 can all be interpreted as parametric variations of this architecture. They differ primarily along two axes: (1) the representational basis (dense semantic embeddings versus sparse lexical vectors), and (2) whether those representations are learned from data or manually engineered. Existing methods based on this bi-encoder architecture vary primarily in their representation format (dense versus sparse), pooling strategies, and training methodologies.
6.3.1 Learned Dense Retrieval
@@
Dense retrieval models typically adopt a standardized dual-encoder architecture built on pre-trained transformer models, most commonly BERT in this context. The standard formulation uses separate BERT encoders for queries and documents, with a layer-normalized linear projection applied to the token representation: $\text{Encoder}(\cdot)=\text{Linear}\big(\text{BERT}(\cdot)\big)$. The encoder weights can be separate or shared between query and document sides.
@@
Relevance scoring in dense retrieval is performed through simple similarity functions, most commonly dot product or cosine similarity between query and document vectors. This design enables efficient ANN search over pre-computed document representations (Johnson et al., 2019), making dense retrieval practical for large-scale collections while maintaining the semantic understanding capabilities of transformer models. The success of this architecture stems from its ability to learn semantic representations that address the vocabulary mismatch problem inherent in traditional sparse retrieval methods (Lee et al., 2019; Karpukhin et al., 2020; Xiong et al., 2020; Reimers & Gurevych, 2019). Dense retrieval models have demonstrated notable effectiveness improvements over BM25 baselines across various tasks including open-domain question answering and web search.
6.3.2 Learned Sparse Retrieval
@@
Learned sparse retrieval (LSR, Figure˜3(c)) employs the same bi-encoder architecture as dense retrieval but produces fundamentally different representations. [^3] We focus on learned sparse retrieval within the bi-encoder formulation, thereby excluding approaches based on learned document expansion (e.g., Nogueira et al. (2019a); Zbib et al. (2019)) and document term reweighting (e.g., DeepCT (Dai & Callan, 2019b)). For a unified treatment of learned sparse retrieval that encompasses these lines of work, see Mallia et al. (2021); Basnet et al. (2024). While sharing the transformer backbone, sparse retrieval models encode queries and documents into high-dimensional sparse vectors whose dimensionality typically matches the vocabulary size of the underlying pre-trained model, often containing tens of thousands of dimensions. Each dimension corresponds to a specific vocabulary term, creating an interpretable representation where non-zero weights indicate term importance (Formal et al., 2021b; a; Nguyen et al., 2023).
@@
Three key architectural constraints distinguish sparse encoders from their dense counterparts. First, sparsity is enforced through explicit regularization techniques, ensuring most term weights remain zero to maintain efficiency (Formal et al., 2021b; Xu et al., 2025b; 2026). Second, all weights must be non-negative to maintain compatibility with traditional inverted index software designed for lexical search systems like Lucene. Third, the high-dimensional vocabulary-aligned vectors enable integration with existing inverted index infrastructure and optimization algorithms (Turtle & Flood, 1995; Broder et al., 2003; Bruch et al., 2024). Notably, inference-free learned sparse retrieval methods such as SPLADE-Doc (Formal et al., 2021a; Shen et al., 2025) eliminates requirement of specialized accelerators such as GPUs, making them highly efficient for inference on multi-core CPU machines.
@@
At a conceptual level, learned sparse retrieval can be viewed as a sophisticated evolution of traditional term weighting schemes, learning context-aware token importance scores from data rather than relying on heuristic formulas (Zamani et al., 2018; Dai & Callan, 2019b; Mallia et al., 2021; Yu et al., 2024a; Xu et al., 2025b; 2026). This approach inherits desirable properties from bag-of-words models such as exact term matching while leveraging the semantic understanding capabilities of pretrained transformers. Notable implementations include SPLADE (Formal et al., 2021a), DeepImpact (Mallia et al., 2021), and uniCOIL (Lin & Ma, 2021), which demonstrate that transformer-based sparse representations can achieve effectiveness comparable to dense retrieval while maintaining the efficiency benefits of inverted indexes.
6.4 Bridging the Gap: Advanced Interaction and Hybrid Models
@@
6.4.1 Multi-Vector Representations
@@
To re-introduce query-document interaction without the full cost of a cross-encoder, multi-vector models represent queries and documents using multiple vectors. Poly-Encoder (Humeau et al., 2020) computes a fixed number of vectors per query and aggregates them with softmax attention over document vectors. ME-BERT (Luan et al., 2021) represents documents with $m$ vectors and uses the maximum similarity between any query and document vector to estimate relevance.
In line with this idea, ColBERT (Khattab & Zaharia, 2020; Santhanam et al., 2022; Hofstätter et al., 2022) represents each token in the query and document as a contextualized vector. It then performs a “late interaction” step where each query vector is compared against all document vectors via a MaxSim operator, and the final score is the sum of these maximum similarities. This late interaction scheme (Figure˜3(d)) allows ColBERT for end-to-end training to achieve strong performance while still achieving efficient retrieval through a dedicated index structure. On the other hand, it also leads to drastically increased index size, which has been the focus in later studies (Santhanam et al., 2022; Hofstätter et al., 2022, inter alia).
We should also note that multi-vector retrieval can be viewed as a special case of dense retrieval where the learned feature representation is a matrix of size $n\times h$, with $n$ vectors of hidden dimension $h$. This matrix can be conceptually flattened into a single dense vector, showing its connection to the vanilla single-vector retrieval. The key difference lies not in the representation itself but in the richer relevance estimation strategy: instead of applying a simple linear relevance like dot product, models like ColBERT aggregate fine-grained token-level interactions to compute a relevance score.
6.4.2 Hybrid Retrieval
@@
Another direction combines the strengths of different retrieval systems. A simple yet effective approach is ranklist fusion (e.g., Reciprocal Rank Fusion, Cormack et al., 2009), which merges ranked lists from sparse (e.g., BM25) and dense retrievers post-retrieval without architectural changes. More integrated models combine signals at a deeper level. COIL (Gao et al., 2021a) enhances traditional bag-of-words retrieval with semantic embeddings from a BERT encoder. uniCOIL (Lin & Ma, 2021) simplifies this by reducing the semantic embedding to a single dimension, effectively learning a term weight akin to LSR models like SPLADE (Formal et al., 2021a; b). A few works fall into the intersection of learned sparse retrieval and multi-vector representations. For example, SLIM (Li et al., 2023c) first maps each contextualized token vector to a sparse, high-dimensional lexical space before performing late interaction between these sparse token embeddings. SPLATE (Formal et al., 2024) takes an alternative approach to first encode contextualized token vectors, then map these token vectors to a sparse vocabulary space with a partially learned SPLADE module. Both models achieve performance improvement compared to learned sparse retrieval baselines such as SPLADE.
Table 3: Summary of IR model architecture for passage retrieval and passage ranking based on pre-trained transformers. Dense Retrieval and LSR denote learned dense retrieval and learned sparse retrieval, respectively. DeepCT (Dai & Callan, 2019b) is trained without labeled training set. Contrastive Learning and in-batch negatives means listwise loss function is used. SentenceBERT (Reimers & Gurevych, 2019) is originally designed for the symmetrical sentence similarity tasks, but is quickly expanded to asymmetrical retrieval tasks.
@@
CEDR (MacAvaney et al., 2019b) Reranking Cross-encoder BERT Contrastive Learning
@@
SentenceBERT (Reimers & Gurevych, 2019) Dense Retrieval Bi-encoder BERT Triplet
@@
SLIM (Li et al., 2023c) LSR + Multi-vector Bi-encoder BERT Contrastive Learning
SLIM++ (Li et al., 2023c) LSR + Multi-vector Bi-encoder BERT Distillation
@@
6.5 Orthogonal Improvements and Analysis
Beyond architectural innovations, performance can be enhanced through improvements to the underlying models and a deeper analysis of their behavior. We show a list of models and their corresponding architectures in Table˜3, a majority of which use BERT (Devlin et al., 2019) as the backbone, with exceptions using DistilBERT (Sanh et al., 2019), RoBERTa (Liu et al., 2019), and T5 (Raffel et al., 2020; Sanh et al., 2022; Mo et al., 2023; Chung et al., 2024).
@@ 6.5 Orthogonal Improvements and Analysis
Instead of modifying the retrieval architecture, this line of work enhances the backbone language model itself through domain adaptation or continued pre-training, a proven strategy in NLP (Howard & Ruder, 2018; Gururangan et al., 2020). For instance, Lee et al. (2019) pre-train BERT with an Inverse-Cloze Task (Taylor, 1953) for better text representations. Condenser (Gao & Callan, 2021) proposes a dedicated pre-training architecture to “condense” text representations into the [CLS] token. COCO-DR (Yu et al., 2022c) extends Condenser by using a technique named Distributionally Robust Optimization to mitigate distribution shifts in dense retrieval. A line of works have also explored other pre-training objectives such as masked auto-encoders (Xiao et al., 2022; Wu et al., 2023b) and bag-of-words prediction (Ma et al., 2024a). Recent works (Wang et al., 2023a; Nussbaum et al., 2025; Yu et al., 2024b, inter alia) have employed a “middle-stage” training on large-scale unlabeled text pairs, between pretrained encoder models and supervised fine-tuning on labeled text pairs, and have demonstrated the corresponding performance improvement compared to the traditional two-stage pipeline. We refer readers to the original papers for details.
@@
A few works have attempted to interpret what transformer-based models learn, i.e., mechanistic interpretability (Elhage et al., 2021; Saphra & Wiegreffe, 2024). MacAvaney et al. (2022) showed that neural models rely less on exact match signals and instead encode rich semantic information. Ram et al. (2023a) connected dense and sparse retrieval by projecting a dense retriever’s intermediate representations into the vocabulary space. Separately, other work focuses on designing systems that provide model-agnostic explanations (Rahimi et al., 2021; Yu et al., 2022b; Xu et al., 2024c) to satisfy desiderata like faithfulness (Jacovi & Goldberg, 2020; Xu et al., 2023). As IR systems become integral to other applied ML domains, we believe it is important to study and design interpretable, truthful, and trustworthy IR models.
The architectural innovations discussed in this section highlight a mature research field dedicated to harnessing pre-trained transformers for information retrieval. The central theme has been the architectural tradeoff between interaction depth and computational cost, giving rise to a spectrum of models from highly effective cross-encoder rerankers to efficient bi-encoder retrievers. By developing advanced representations — whether dense, sparse, or multi-vector — and hybridizing these approaches, the community has pushed the boundaries of the classic “retrieve-then-rank” paradigm. However, these models still primarily function as specialized components for representation and scoring. The next wave of innovation emerged from models capable not only of understanding text but also of generating it, marking the beginning of the era of Large Language Models.